Give the systematic name for the following disacchGive the systematic name for the following disaccharide by selecting the appropriate terms and filling in the blank number field.
Answer #1
Concepts and reason

The concept used to solve this question is to write the systematic name of the given disaccharide. Disaccharides are carbohydrates which are formed by the combination of two monosaccharide (sugar) units. The two monosaccharide units are bonded by a glycoside linkage (C-O-C bond). To write the systematic name of the given disaccharide, the name of the two monomer sugar units and position of glycoside linkage should be identified.

Fundamentals

The carbohydrates which cannot be hydrolyzed further to smaller compounds are known as monosaccharides. Therefore, these are also called as simple sugars. For example, (+)-glucose, fructose, etc.

Disaccharides: Combination of monosaccharide unit gives more complex carbohydrates. Disaccharides are carbohydrates which are formed by the combination of two monosaccharide (sugar) units. The most common examples of disaccharides are maltose, lactose, and sucrose.

The combination of two monomeric sugar units takes place by the elimination of water molecule. During the elimination a new C-O-C bond is formed between two rings which is called glycoside linkage. In disaccharides, the individual sugar units are in ring form, pyranose or furanose form.

The glycoside linkage in the disaccharide is formed between the anomeric carbon of one monomer unit and hydroxy group of another monomer unit.

Nomenclature of disaccharides:

The systematic name of disaccharides is very lengthy but explains the bonding between two monomer units. The following rules can be followed to write the systematic name.

1) First identify the names of monosaccharide units present in the disaccharide molecule.

2) Identify which form of monosaccharide unit is present whether it is in the α\alpha -form or in the β\beta -form. They differ in the orientation of hydroxyl group and hydrogen at the anomeric carbon atom. If the hydroxy group is below the pyranose ring then it is α\alpha -form and if the hydroxy group is below the pyranose ring then it is β\beta -form. For example, the α\alpha - and β\beta -forms of glucose are shown below.

НОН
НОН
IL HO
но
НО
HOOH
Н
Ho H oH
Но то
н
HO
ПОН
Above the plane
abov
a-form
B-form
Below the plane

3) Identify the link of the glycoside linkage between the two monosaccharide units. For example if the anomeric carbon of one monomer unit is bonded to the hydroxy group of C4 carbon of other monomer unit then the linkage is called as 141 \to 4 linkage.

The given disaccharide molecule is shown below.

НО
ТНО НЯ
ОН
НО
НОН
нхон он
|
0
н
1
но
|
н
н
о н
т
НО
н

The straight chain structures of D-glucose and D-galactose are shown below.

CHO
н-он
Нонн
нон
ньон
СН,ОН
D-Glucose
CHO
НЕОН
но н
Нонн
НЕОН
СН,ОН
D-Galactose

The pyranose structures of D-glucose and D-galactose including both α\alpha - and β\beta -forms are shown below.

НОН
НОН
ОН
.
но
НО ОН
но
HỌ
Ho
Lo
но
H
HO
НОМ
НО
НО
ОН
но
НЕ
НО
н \
-ОН
н Онон
Н ОНЫ
a-form
B-form
Pyronose structures of D-g

The only difference in the structure of D-glucose and D-galactose that the hydroxy group at C4 carbon in D-glucose is present below the plane whereas the hydroxy group at C4 carbon in D-galactose is present above the plane.

By keenly observing at the above structures, α\alpha -D-glucopyranose unit is present as two monomeric units in the given disaccharide molecule.

It is observed that the α\alpha -D-glucopyranose units are linked by the glycoside in the given disaccharide molecule.

Next, identify the position of linkage in the given disaccharide molecule by numbering the disaccharide as follows:

но но
н
н
ТОН
НО-
O
HO
WH
H
Но
5 он
н
н
н
НО

After numbering, it is observed that the glycoside linkage is formed between the anomeric carbon of one monomer unit (C1) and hydroxy group of C4 carbon of other monomer unit, thus the linkage is called as 141 \to 4 linkage.

Therefore, the systematic name of the given disaccharide can be written as 4-O- α\alpha -D-glucopyranosyl- α\alpha -D-glucopyranose.

Ans:

The appropriate terms for the systematic name of the given disaccharide from the list are α\alpha -D-glucopyranosyl and α\alpha -D-glucopyranose, the correct answer for the blank is 4.